# Large Language Models as Neurolinguistic Subjects: Identifying Internal Representations for Form and Meaning

### v1

@@ Abstract
This study investigates the linguistic understanding of Large Language Models (LLMs) regarding signifier (form) and signified (meaning) by distinguishing two LLM assessment paradigms: psycholinguistic and neurolinguistic. Traditional psycholinguistic evaluations often reflect statistical biases that may misrepresent LLMs’ true linguistic capabilities. We introduce a neurolinguistic approach, utilizing a novel method that combines minimal pair and diagnostic probing to analyze activation patterns across model layers. This method allows for a detailed examination of how LLMs represent form and meaning, and whether these representations are consistent across languages. Our contributions are three-fold: (1) We compare neurolinguistic and psycholinguistic methods, revealing distinct patterns in LLM assessment; (2) We demonstrate that LLMs exhibit higher competence in form compared to meaning, with the latter largely correlated to the former; (3) We present new conceptual minimal pair datasets for Chinese (COMPS-ZH) and German (COMPS-DE), complementing existing English datasets.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable reasoning, linguistic, arithmetic, and other cognitive abilities. The advent of LLMs has reignited cross-disciplinary discussions about what sorts of behavior are “intelligence”, even if the intelligence exhibited by LLMs may differ from human intelligence Sejnowski (2023). LLMs have drawn the attention of researchers from various fields, including linguistics, cognitive science, computer science and neuroscience, who investigate how LLMs develop and exhibit these capabilities.
Figure 1: Illustration of LLMs prcocessing the same signified (meaning) across different signifiers (forms).
Figure 2: Psycholinguistic vs. Neurolinguistic Paradigm. Both direct probability measurement and metalinguistic prompting can be considered as psycholinguistic methods, while minimal pair probing He et al. (2024) and other diagnostic probing are neurolinguistic.
There is currently a heated debate about whether LLMs understand human language or whether their performance is simply the product of complex statistical relationships Mitchell and Krakauer (2023). Some assert that LLMs have demonstrated the potential for artificial general intelligence Bubeck et al. (2023). According to Manning (2022), optimism grows that emerging knowledge-imbued systems have exhibited a degree of general intelligence. Others remain skeptical Bender and Koller (2020). Browning and LeCun (2022) argue that LLMs are limited to a superficial understanding. Bender et al. (2021) suggests that LLMs cannot possess true understanding because they lack mental models of the world and experiential knowledge.
A central aspect of this debate concerns the nature of LLMs’ linguistic representations. Using the semiotic framework of language proposed by De Saussure (1989), which distinguishes between the signifier (form) and the signified (meaning), we can inquire into the extent to which LLMs comprehend the form and meaning, and how do form and meaning intertwist with each other. Is LLM’s understanding of language meaning merely a statistical outcome based on their grasp of language form? When different languages express a shared concept with distinct forms, do LLMs create similar representations for these variations? How can we better understand the representations of form and meaning in these systems that support the observed patterns of performance?
@@
Drawing inspirations from cognitive study on human language processing, we propose that assessment of LLMs can be divided into two primary paradigms: psycholinguistic and neurolinguistic. As illustrated in Figure 2, psycholinguistic paradigm measures the model’s output probabilities, directly reflecting the model’s behavior and performance. Neurolinguistic paradigm delves into the internal representations of LLMs.
While prior studies have treated language models as psycholinguistic subjects Futrell et al. (2019), this research is the first to propose viewing LLMs as neurolinguistic subjects as well.
When treating LLMs as psycholinguistic subjects, their responses often lean towards signifiers (form) due to their reliance on statistical correlations. This enables LLMs to produce structurally coherent language but not necessarily semantically accurate responses, as their “understanding” is shaped by statistical patterns rather than true conceptual processing Harnad (1990); Bender and Koller (2020); Nie et al. (2024), which makes psycholinguistic evaluation results might not accurately reflect the true linguistic understanding of LLMs.
In contrast, examining LLMs as neurolinguistic subjects focuses on internal representations, which may transcend surface-level statistical biases. To achieve neurolinguistic assessing, we adapted the decoding probing method presented by He et al. (2024), which we refer to as “minimal pair probing” to distinguish it from conventional diagnostic probing Belinkov and Glass (2019); Belinkov (2022). Using this method, we can examine how LLMs represent both form and meaning in a fine-grained, layer-by-layer manner.
In order to address questions about whether LLMs maintain consistent underlying representations of the same concept when form changes across multiple languages, we also create a multilingual minimal pair dataset (COMPS-ZH for Chinese and COMPS-DE for German).
By evaluating LLMs in both psycholinguistic and neurolinguistic paradigm, our results reveal:1) Psycholinguistic and neurolinguistic results reveal very different patterns, suggesting both paradigms are necessary for a comprehensive understanding of LLMs. 2) LLMs demonstrate easier, earlier and better competence in language form than meaning. 3) As linguistic form varies across different languages, the LLMs’ understanding of meaning appears to shift accordingly with the meaning competence being linearly correlated to form. Results together suggest that signifier and signified in LLMs might not be independent to each other. Maintaining a conceptual representation might rely on statistical correlations based on linguistic form.
@@ 2.2 In LLM Assessment Research
A language model is considered to perform correctly on this task if it assigns a higher probability to the acceptable sentence compared to the unacceptable one Marvin and Linzen (2018). Researchers created syntactic, semantic/conceptual, and discourse inference tasks for the minimal pair method. They provide more precise insights into the abilities of LLMs compared to metalinguistic prompting Futrell et al. (2019); Gauthier et al. (2020); Hu et al. (2020); Warstadt et al. (2020); Beyer et al. (2021); Misra et al. (2023); Kauf et al. (2023).
Through either metalinguistic judgement or direct probability measurement methods, these tasks essentially treat LLMs as psycholinguistic subjects Futrell et al. (2019). This research paradigm resembles cognitive psychology by having LLMs perform tasks, such as cloze and question answering, and then evaluating their performance without examining the internal representations. Information about the inner workings of a model is inferred either from its output or from the probabilities it assigns to different possible outputs. The internal states of the LLM (i.e. its intermediate layers) are not examined.
@@
Another line of research focuses on studying the internal representations, emphasizing a neurolinguistic approach to understanding LLMs. Essentially, previous diagnostic probing methods in evaluating language models can be considered as neurolinguistic paradigms as they examine the internal states of LMs Belinkov and Glass (2019); Belinkov (2022), while the term ‘neurolinguistic’ hasn’t been applied to the field before. Diagnostic probing involves training a classifier to predict linguistic properties from the hidden states of LMs. Following this paradigm, researchers decode syntactic, semantic, morphological, and other linguistic properties from the hidden states of LMs Köhn (2015); Gupta et al. (2015); Shi et al. (2016); Tenney et al. (2019); Manning et al. (2020).
@@ 3 Minimal Pair Probing = Minimal Pair + Diagnostic Probing
While prior neurolinguistic approaches have explored internal representations, they often employed coarse-grained datasets and primarily focused on decoding linguistic labels from embeddings, providing a general perspective on the linguistic features encoded in LMs. In contrast, the minimal pair probing method presented by He et al. (2024) integrates minimal pair design with diagnostic probing. This combination leverages the granularity of minimal pair design and the layer-wise insights of diagnostic probing, thereby enabling a more detailed analysis of internal patterns for form and meaning. We adopt the minimal pair decoding method as the neurolinguistic paradigm in our work.
Specifically, given an LLM $f:x_{0,1,...,i}\rightarrow x_{i+1}$ trained on dataset $\mathcal{D}_{O}$, we can extract the hidden state representations $f_{l}(S)$ of the $l$-th layer of stimuli $S$. Given a minimal pair dataset $\mathcal{D}_{m}$ = $\{(S_{+}^{i},S_{-}^{i}),(z_{+}^{i},z_{-}^{i})\}$ with each sentence $S$ has a label $z$, we have internal representation $f_{l}(S_{+}^{i})$ and $f_{l}(S_{-}^{i})$ for each sentence. A minimal probing classifier $g:f_{l}(S)\rightarrow\hat{z}$ is trained and evaluated on $\mathcal{D}_{m}$, with grammatical/conceptual performance measure $\mathrm{Perf}(f,\mathcal{D}_{O},g,\mathcal{D}_{m})$. Note that what we focus on is the performance of the LLM $f$ itself, i.e., $\mathrm{Perf}(f,\mathcal{D}_{O})$, instead of the capability of the supervised classifier $g$. To eliminate the influence of $g$, we assign a random vector $r$ to each sentence to replace the hidden states $f_{l}(S)$, and calculate the performance of the classifier $g$ on the random vector $r$ as $\mathrm{Perf}(g,\mathcal{D}_{m})$. This “random” baseline evaluates the predictability of the probing classifier itself. Finally, the performance of the LLM $f$ is defined as follows:
4 Experiment Setup
4.1 Datasets
We use minimal pair probing for English, Chinese, and German to assess grammaticality (form) and conceptuality (meaning). Table 1 presents the overall dataset information used in our experiments. Further dataset descriptions are in Appendix B.
4.2 Models
Resource Level Llama2 Llama3 Qwen
English High High High
Chinese Mid Mid High
German Low Low Low
Llama2 and Llama3
QWen
@@ 4.2 Models
Table 3: Prompt examples for baseline methods. Region where we measure probability is marked in color. Correct sentences and answers are in blue; incorrect in red.
4.3 Setup for Psycholinguistic Paradigms
@@
Direct probability measurement calculates the probability of entire sentences based on model logits. Accuracy is determined by whether the model assigns a higher probability to the grammatically or conceptually correct sentence within the minimal pair.
@@
Metalinguistic prompting involves explicitly asking a question or specifying a task that requires a judgement about a linguistic expression. Following Hu and Levy (2023), we use a single prompt each minimal pair to present both sentences at once. For form tasks, we assign an identifier (1 or 2) to each sentence in the pair, present a multiple-choice question comparing both sentences, and compare the probabilities assigned by the model to each answer option, “1” or “2”. For meaning tasks, we reformulate the property into a question and compare the probabilities of acceptable and unacceptable concepts as sentence continuations. Table 3 presents the prompts used in the experiments.
4.4 Setup for Minimal Pair Probing
@@ 5.1 Psycholinguistic vs Neurolinguistic
Figure 3: Psycholinguistic vs. Neurolinguistic Results. Direct refers to direct probability measurement; meta refers to metalinguistic prompting and neuro refers to minimal pair probing, which is a neurolinguistic paradigm.
@@
In order to compare psycholinguistic and neurolinguistic paradigms, we use the last layer’s performance in the neuro method  [^1] We refer to minimal pair probing as the neuro method for simplicity in the results section., as both direct probability measurement and metalinguistic prompting rely on the last layer of LLMs.
Differences among Direct, Meta, and Neuro results
Figure 3 shows distinct results for Direct, Meta, and Neuro, with a general performance order of Neuro > Direct > Meta. This suggests that relying on a single method may not give a comprehensive assessment of language models, and it is best to employ multiple methods for evaluation.
Neuro is similar to Direct on meaning but not on form.
The Neuro method consistently outperforms others across models when assessing form, often nearing 1.0, while Direct and Meta show more variability and generally lower performance. The Neuro method performs similarly to Direct in meaning evaluation, yet remains significantly higher than Meta.
Meta method performs consistently poorly.
This serves as a caution for the current trend of instruction-based prompting, suggesting that Meta’s limitations could reflect issues with instruction prompting’s effectiveness.
Instruction tuning influences LLMs’ Meta performance on form but not much on meaning.
Chat version models’ meta performance on form varies a lot compared to their base versions, but meaning understanding keeps stable.
5.2 Detailed Neurolinguistic Results within English
LLMs incrementally encode grammar and concepts.
Figure 4: Detailed English minimal pair probing results. The pink, yellow, and blue curves represent three grammatical levels: morphology, the semantics-syntax interface, and syntax, respectively. The red curves denote conceptual tasks.
To investigate how LLMs encode form and meaning in a fine-grained and layer-by-layer manner, we split form back into 12 original grammatical tasks and meaning into 4 original conceptual tasks. Figure 4 indicates that Llama-2 rapidly learns grammatical and conceptual information in the first layers incrementally. Performance improves across layers and stabilizes at some point. Similar results for all 6 models can be found in Figure 14 of Appendix D.
Figure 4 shows result curves for 16 specific linguistic tasks. 12 curves (with pink, yellow, and blue colors) visualize the results of grammatical tasks, while the 4 red curves represent the results of the conceptual tasks.
LLMs encode grammatical features better than conceptual features.
Figure 5: English minimal pair probing results for form and meaning.
@@ 5.2 Detailed Neurolinguistic Results within English
The results in Figure 5 show that the performance scores for conceptual understanding (red curves) are significantly lower than those for grammatical understanding (blue curves). This pattern is consistent across all six models, suggesting a universal characteristic of LLMs.
LLMs encode meaning after form.
Figure 6: Feature learning saturation (95% of peak) and maximum layer.
@@
We compute the feature learning saturation and maximum layers for all 12 grammatical tasks and 4 conceptual tasks, averaging them to represent form and meaning, respectively. From Figure 6, we can see that the saturation and maximum layers for meaning are generally higher than those for form across all six models. This suggests that LLMs stabilize their encoding of grammatical features before conceptual features, indicating that LLMs learn the form of language first and then the meaning of language.
Instruction tuning won’t change internal linguistic representation much.
Figure 7: Performance difference between the base and the instruction versions of LLMs. The shaded areas indicate standard deviation where $\text{std}_{\text{diff}}=\sqrt{\text{std}_{\text{base}}^{2}+\text{std}_{\text{chat}}^{2}}$.
From Figure 7, instruction tuning does not significantly alter the internal linguistic representations of the models. The performance differences for both form and meaning remain close to zero throughout all layers, suggesting that the overall linguistic encoding remains stable despite the tuning.
5.3 Meaning Encoding across Different Forms

### v2

# Large Language Models as Neurolinguistic Subjects: Discrepancy in Performance and Competence for Form and Meaning

@@ Abstract
This study investigates the linguistic understanding of Large Language Models (LLMs) regarding signifier (form) and signified (meaning) by distinguishing two LLM assessment paradigms: psycholinguistic and neurolinguistic. Traditional psycholinguistic evaluations often reflect statistical rules that may not accurately represent LLMs’ true linguistic competence. We introduce a neurolinguistic approach, utilizing a novel method that combines minimal pair and diagnostic probing to analyze activation patterns across model layers. This method allows for a detailed examination of how LLMs represent form and meaning, and whether these representations are consistent across languages. We found: (1) Psycholinguistic and neurolinguistic methods reveal that language performance and competence are distinct; (2) Direct probability measurement may not accurately assess linguistic competence; (3) Instruction tuning won’t change much competence but improve performance; (4) LLMs exhibit higher competence and performance in form compared to meaning. Additionally, we introduce new conceptual minimal pair datasets for Chinese (COMPS-ZH) and German (COMPS-DE), complementing existing English datasets.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated remarkable reasoning, linguistic, arithmetic, and other cognitive abilities. The advent of LLMs has reignited cross-disciplinary discussions about what sorts of behavior are “intelligence”, even if the intelligence exhibited by LLMs may differ from human intelligence Sejnowski (2023). LLMs have drawn the attention of researchers from various fields, including linguistics, cognitive science, computer science, and neuroscience, who investigate how LLMs develop and exhibit these capabilities.
Figure 1: Left: Research Question. Right: Psycholinguistic vs. Neurolinguistic Paradigm. Both direct probability measurement and metalinguistic prompting can be considered psycholinguistic methods, while minimal pair probing He et al. (2024) and other diagnostic probing are neurolinguistic.
There is currently a heated debate about whether LLMs understand human language or whether their performance is simply the product of complex statistical relationships Mitchell and Krakauer (2023). A central aspect of this debate concerns the nature of LLMs’ linguistic representations. Using the semiotic framework of language proposed by De Saussure (1989), which distinguishes between the signifier (form) and the signified (meaning), we can inquire into the extent to which LLMs comprehend the form and meaning, and how form and meaning intertwist with each other. Is LLM’s understanding of language meaning merely a statistical outcome based on their grasp of language form? When different languages express a shared concept with distinct forms, do LLMs create similar representations for these variations? How can we better understand the representations of form and meaning in these systems that support the observed patterns of performance?
@@
Drawing inspirations from the cognitive study on human language processing, we propose that the assessment of LLMs can be divided into two primary paradigms: psycholinguistic and neurolinguistic. As illustrated in Figure 1, the psycholinguistic paradigm measures the model’s output probabilities, directly reflecting the model’s behavior and performance. The neurolinguistic paradigm delves into the internal representations of LLMs.
When treating LLMs as psycholinguistic subjects, their responses may leverage their grasp of form, relying on statistical correlations, to create an illusion of understanding meaning. This enables LLMs to produce structurally coherent but not necessarily semantically accurate responses, as their “understanding” is shaped by patterns rather than true conceptual processing Harnad (1990); Bender and Koller (2020); Nie et al. (2024). Consequently, psycholinguistic evaluations tend to reflect performance rather than competence, as they assess external outputs that may not fully capture the underlying linguistic knowledge encoded within the model. This mismatch suggests that psycholinguistic evaluation results might not accurately represent the true linguistic competence of LLMs.
In contrast, examining LLMs as neurolinguistic subjects focuses on internal representations, providing a more direct assessment of competence by moving beyond surface-level biases Firestone (2020). To achieve this, we adapted the decoding probing method by He et al. (2024), referred to as “minimal pair probing”, to analyze how LLMs encode form and meaning across layers. This approach allows for a finer distinction between performance and competence, revealing insights that psycholinguistic methods might overlook.
In order to address questions about whether LLMs maintain consistent underlying representations of the same concept when the form changes across multiple languages, we also create a multilingual minimal pair dataset (COMPS-ZH for Chinese and COMPS-DE for German).
By evaluating LLMs in both psycholinguistic and neurolinguistic paradigms, we found: 1) Psycholinguistic and neurolinguistic results reveal very different patterns, suggesting both paradigms are necessary for a comprehensive understanding of LLMs. 2) Though more intrinsic than metalinguistic prompting, direct probability measurement may still not accurately assess linguistic competence, as it remains influenced by statistical patterns. 3) LLMs acquire competence in linguistic form more easily, earlier, and with greater accuracy than in meaning. 4) As linguistic form varies across languages, LLMs’ understanding of the same concept shifts accordingly, with meaning competence linearly correlated to form. This suggests that signifier and signified in LLMs may not be independent, and maintaining conceptual representations likely depends on statistical correlations with form.
@@ 2.2 In LLM Assessment Research
A language model is considered to perform correctly on this task if it assigns a higher probability to the acceptable sentence compared to the unacceptable one Marvin and Linzen (2018). Researchers have created syntactic, semantic/conceptual, and discourse inference tasks for the minimal pair method. They provide more precise insights into the abilities of LLMs compared to metalinguistic prompting Futrell et al. (2019); Gauthier et al. (2020); Hu et al. (2020); Warstadt et al. (2020); Beyer et al. (2021); Misra et al. (2023); Kauf et al. (2023).
Through either metalinguistic judgement or direct probability measurement methods, these tasks essentially treat LLMs as psycholinguistic subjects Futrell et al. (2019). This research paradigm resembles cognitive psychology by having LLMs perform tasks, such as cloze and question answering, and then evaluating their performance without examining the internal representations, in a manner similar to how subjects participate in psychological experiments. Information about the inner workings of a model is inferred either from its output or from the probabilities it assigns to different possible outputs. The internal states of the LLM (i.e. its intermediate layers) are not examined.
@@
Another line of research focuses on studying the internal representations, emphasizing a neurolinguistic approach to understanding LLMs. Essentially, diagnostic probing methods in evaluating language models can be considered as neurolinguistic paradigms as they examine the internal states of LMs Belinkov and Glass (2019); Belinkov (2022), while the term ‘neurolinguistic’ hasn’t been applied to the field before. Diagnostic probing involves training a classifier to predict linguistic properties from the hidden states of LMs. Following this paradigm, researchers decode syntactic, semantic, morphological, and other linguistic properties from the hidden states of LMs Köhn (2015); Gupta et al. (2015); Shi et al. (2016); Tenney et al. (2019); Manning et al. (2020).
@@ 3 Minimal Pair Probing = Minimal Pair + Diagnostic Probing
Table 2: Prompt examples for baseline methods. The region where we measure probability is marked in color. Correct sentences and answers are in blue; incorrect in red.
While prior neurolinguistic approaches have explored internal representations, they often employed coarse-grained datasets and primarily focused on decoding linguistic labels from embeddings, providing a general perspective on the linguistic features encoded in LMs. In contrast, the minimal pair probing method presented by He et al. (2024) integrates minimal pair design with diagnostic probing. This combination leverages the granularity of minimal pair design and the layer-wise insights of diagnostic probing, thereby enabling a more detailed analysis of internal patterns for form and meaning. We adopt minimal pair decoding as the neurolinguistic paradigm in our work.
Specifically, given an LLM $f:x_{0,1,...,i}\rightarrow x_{i+1}$ trained on dataset $\mathcal{D}_{O}$, we can extract the hidden state representations $f_{l}(S)$ of the $l$-th layer of stimuli $S$. Given a minimal pair dataset $\mathcal{D}_{m}$ = $\{(S_{+}^{i},S_{-}^{i}),(z_{+}^{i},z_{-}^{i})\}$ with each sentence $S$ has a label $z$, we have internal representation $f_{l}(S_{+}^{i})$ and $f_{l}(S_{-}^{i})$ for each sentence. A minimal probing classifier $g:f_{l}(S)\rightarrow\hat{z}$ is trained and evaluated on $\mathcal{D}_{m}$, with grammatical/conceptual performance measure $\mathrm{Perf}(f,\mathcal{D}_{O},g,\mathcal{D}_{m})$. Note that our focus is the performance of the LLM $f$ itself, i.e., $\mathrm{Perf}(f,\mathcal{D}_{O})$, instead of the capability of the supervised classifier $g$. To eliminate the influence of $g$, we set up a “random” baseline, where we assign a random vector $r$ to each sentence to replace the hidden states $f_{l}(S)$ and calculate the performance of the classifier $g$ on the random vector $r$ as $\mathrm{Perf}(g,\mathcal{D}_{m})$, aiming to evaluate the inherent predictability of the probing classifier. Finally, the performance of the LLM $f$ is defined as follows:
4 Experiment Setup
4.1 Datasets and Models
We use minimal pair probing for English, Chinese, and German to assess grammaticality (form) and conceptuality (meaning). Table 1 presents the overall dataset information used in our experiments. We use Llama2-7B, Llama3-8B, and Qwen-7B models in both base and chat versions. Further dataset and model descriptions are in Appendix B and C.
4.2 Setup for Psycholinguistic Analysis
@@
Direct probability measurement calculates the probability of a sentence based on model logits. Accuracy is determined by whether the model assigns a higher probability to the grammatically or conceptually correct sentence within the minimal pair.
@@
Metalinguistic prompting involves explicitly asking a question or specifying a task that requires a judgement about a linguistic expression. Following Hu and Levy (2023), we use one prompt for a minimal pair to present both sentences at once. For form tasks, we assign an identifier (1 or 2) to each sentence in the pair, present a multiple-choice question comparing both sentences and compare the probabilities assigned by the model to each answer option, “1” or “2”. For meaning tasks, we reformulate the property into a question and compare the probabilities of acceptable and unacceptable concepts as sentence continuations. Table 2 presents the prompts used in the experiments.
Figure 2: Psycholinguistic (meta and direct) and neurolinguistic performance across models and linguistic tasks. The x-axis represents different models and conditions (base and chat), while the y-axis categorizes linguistic tasks based on structural (syntax, morphology, syntax-semantics interface) and conceptual (meaning) levels.
4.3 Setup for Neurolinguistic Analysis
@@
Unsupervised Analysis
We use t-SNE to visualize the sentence embedding of Llama2-7B for English form tasks. We employ PCA to reduce the dimensionality of the sentence embedding to 50 before applying t-SNE.
@@ 5 Results
Figure 3: (a) Averaged psycholinguistic (meta and direct) and neurolinguistic results across models and tasks. t-tests were conducted on the original (pre-averaging) results between base and chat models, with p-values annotated. (b) Psycholinguistic and neurolinguistic performance for form (morphology, syntax-semantics interface, and syntax) and meaning (concept). (c) Competence and performance gap drops after instruction tuning.
@@ 5.1 Psycholinguistic vs Neurolinguistic
Figure 2 shows the performance of LLMs across all linguistic tasks. Figure 3-(a) demonstrates the averaged performance of LLMs across models and 4 levels (syntax, morphology, syntax-semantics interfaces, concept). Figure 3-(b) presents the average performance of LLMs across form and meaning tasks for Direct, Meta, and Neuro [^1] We refer to minimal pair probing as Neuro for simplicity. methods. We use the last layer’s performance in the Neuro method when comparing psycho- and neurolinguistic paradigms, as both direct probability measurement and metalinguistic prompting rely on the last layer of LLMs.
Language performance and competence are distinct (Competence $>$ Performance).
Figure 3 shows distinct results between language performance and competence. Moving from Meta → Direct → Neuro, the evaluation focus gradually shifts from language performance (task execution ability) to language competence (the underlying linguistic ability). Within the same task category, Neuro methods consistently yield higher performance than Direct methods, which in turn outperform Meta methods. This indicates that when evaluating pure linguistic competence, LLMs perform well, but their performance drops when assessed in a task-based setting.
Tasks that emphasize language performance become more difficult, even if their language competence is high.
For example, in the Neuro setting, performance on Syntax tasks reaches 97%, while in the Meta setting, it drops to 56.1%, showing a significant gap. This suggests that even when an LLM has strong competence in a given task, its performance can significantly decline when assessed under performance-oriented evaluation.
Direct probability measurement might not be true competence assessment.
As the Neuro method measures the internal representations of LLMs directly, it could serve as a reliable ground truth for estimating linguistic competence. Direct probability measurement falls short of achieving this ground truth in form assessment (especially for syntax and syntax-semantics-interface as shown in Figure 3-(b)).
LLMs exhibit stronger mastery of form than meaning, regardless of performance or competence.
As shown in Figure 3-(b), LLMs consistently perform better on form-related tasks than on meaning-related tasks. This trend holds regardless of whether the model is a base or chat version. Crucially, this pattern is evident across all evaluation methods. This indicates that LLMs have a stronger grasp of linguistic form than conceptual meaning, whether assessed through task execution or underlying capability.
Instruction tuning won’t change much competence but improve performance.
Neuro results between the base and chat versions of LLMs reveal that instruction fine-tuning does not significantly alter the language competence of the models (t-test between Neuro-Base vs. Neuro-Chat as shown in Figure 3-(a)). With instruction fine-tuning (chat versions of LLMs), Meta performance on form improves significantly while meaning understanding remains stable. Figure 3-(c) illustrates that after instruction tuning, the competence-performance gap (Neuro-Meta) significantly decreases for form-related tasks, while the change for meaning-related tasks remains relatively small. This indicates that fine-tuning with well-designed instructions helps LLMs improve their performance on form-related tasks, bringing them closer to their underlying competence. However, for meaning-related tasks, instruction tuning does not lead to a fundamental improvement in understanding. This indicates that more optimized information access strategies can enhance the external performance of language models, particularly for form-related tasks.
5.2 Neurolinguistic Analysis [^2] Raw results for English, Chinese, and German can be found in Figure 8, 9 and 10 in the Appendix.
Figure 4: t-SNE visualization of embedding differences between acceptable and unacceptable sentences, with red for syntax, purple for morphology, and yellow for the syntax-semantics interface.
Layer-wise unsupervised dynamics reveal gradual emergence of form features
Figure 4 illustrates the layer-wise differences between embeddings for grammatically correct and incorrect sentences. In early layers, the embedding difference appears scattered and unstructured, but as depth increases, they form clearer clusters, indicating a progressively refined sensitivity to syntactic correctness. By Layer 16 and beyond, distinct clusters emerge corresponding to syntax, morphology, and syntax-semantics interface. The results demonstrate that LLMs encode grammaticality judgments dynamically across layers, progressively structuring linguistic representations. Moreover, the formation of distinct clusters for different linguistic phenomena in the unsupervised analysis provides supporting evidence for subsequent supervised classification.
Gradual decline in encoding performance from structure to meaning.
@@
The results in Figure 5-(c) show that the performance scores for conceptual understanding are significantly lower than those for grammatical understanding. This pattern is consistent across all six models, suggesting a universal characteristic of LLMs. Moreover, as illustrated in Figure 5-(a),(b), the encoding performance progressively declines from more structural tasks to more semantic tasks—spanning syntax, morphology, the syntax-semantic interface, and finally conceptual understanding. This highlights that LLMs encode features less effectively as the tasks shift from structure-focused to meaning-focused.
Figure 5: (a) Neuro probing results for 12 tasks in Llama-2 (4 syntax, 4 morphology, 4 syn.-sem. inter., 4 concept). (b) Averaged performance across four linguistic levels in Llama-2. (c) Averaged performance on form vs. meaning across all six models. (d) Feature learning saturation (95% of peak) and maximum layer. (e) Perf. difference between base and instruction-tuned LLMs. (f) Neuro probing results for Chinese and German. (g) Correlation between meaning competence and form competence.
LLMs encode form earlier than meaning.
@@
We compute the feature learning saturation and maximum layers for all 12 grammatical tasks and 4 conceptual tasks, averaging them to represent form and meaning, respectively. As shown in Figure 5-(d), the saturation and maximum layers for meaning are generally higher than those for form across all six models. This suggests that LLMs stabilize their encoding of grammatical features before conceptual features.
Instruction tuning has minimal impact on the internal linguistic representations.
As Figure 5-(e) shows, performance differences for form and meaning remain near zero across all layers, indicating that instruction tuning minimally impacts internal linguistic representations, consistent with our psycholinguistic vs. neurolinguistic analysis.
5.3 Multilingual analysis